# Large Language Models Are Human-Like Internally

### v1

@@ 6.1 Residual error analysis
Following Oh and Schuler (2023b), we analyze by-token squared residual errors from regression models predicting human data (§ 3.3). We specifically use the largest data of DC (FPGD) and identify tokens where the use of the best internal layer, rather than the last layer, notably reduces errors. [^10] Oh and Schuler (2023b) reported a misalignment between MSEs (residual errors) and log-likelihood scores due to the Euclidean norm penalty adopted in the lme4 package. This did not arise in our analysis with statsmodels. To address this question, we fit a linear regression model to predic the decreases in squared residual errors observed in each data point $w_{t}$ with an LM $\theta$, with by-token linguistic properties as features:
$\mathrm{error\_decrease}(w_{t},\theta)\sim\mathrm{model}(\theta)+\mathrm{length}(w_{t})$
$+\mathrm{freq}(w_{t})+\mathrm{position}(w_{t})+\mathrm{POS}(w_{t})\;\;\mathrm{.}$ (6)
@@
6.2 Contextualization in internal layers
We propose an additional perspective linking our findings to the context sensitivity of humans and LMs during sentence processing. Earlier layers may better model human reading behavior because they are less contextualized, reflecting the human-like tendency to process sentences under working memory constraints.
@@ 6.2 Contextualization in internal layers
We also confirmed the by-layer context-sensitivity of surprisal by analyzing the correlation between (i) intermediate layers’ surprisal and less-contextualized bigram surprisal [^11] Bigram LM is trained on OpenWebText Gokaslan and Cohen (2019) with KenLM toolkit Heafield (2011) and (ii) intermediate layers’ surprisal and more contextualized surprisal from the LLM with the lowest PPL (OPT-66B). Figure 5 shows the above two types of correlations for each model’s layer with the x-axis as the relative layer depth. This shows that earlier layers correlate more strongly with bigram surprisal (Pearson correlation coefficient $r$ for this relationship is --0.92), while later layers align with more accurate, well-contextualized surprisal ($r$=0.95). These findings reinforce the idea that earlier layers exhibit limited context sensitivity, while later layers are more contextualized and better suited for modeling data like MAZE, which demands higher contextualization.
Figure 5: The markers correspond to all the internal layers of our targeted LMs, which are sorted by relative layer depth (x-axis). Two types of scores (y-axis) are plotted: (i) Pearson correlation coefficient between each layer’s surprisal vs. less-contextualized bigram surprisal (blue), and (ii) each layer’s surprisal vs. well-contextualized LLM surprisal (red). We used tuned-lens results.
@@ 7 Conclusions
Recent cognitive modeling studies have demonstrated the worse fit of surprisal from larger LMs to human reading time. In this paper, we argue that these negative results stem from the exclusive focus on the final layers of LMs. Instead, we observe that those from internal layers comparably or even better fit with human behavior and neurophysiology data, suggesting that the cognitive plausibility of larger LMs has been underestimated. Furthermore, different human measurements align with different layers, implying an intriguing parallel between temporal dynamics in human sentence processing and LM internal layers.
@@
The experiments can be extended to more types of human measures, such as first fixation time, go-past reading time, eye regressions, total fixation time, EEG data other than the N400 potential Federmeier et al. (2007) as well as FMRI data Shain et al. (2022). As a first step, we have begun with the measures of SPR, FPGD, N400, and MAZE, as they are typically used in cognitive modeling.
Recent studies have raised several issues, orthogonal to our study, on LM-based cognitive modeling, for example, on tokenizations Nair and Resnik (2023); Giulianelli et al. (2024a); Oh and Schuler (2024), LM training scenario Oh and Schuler (2023a), as well as more refined indicators of word predictability Pimentel et al. (2023); Giulianelli et al. (2024b); Opedal et al. (2024); Meister et al. (2024). More critically, the reliance solely on surprisal obtained from (not necessarily cognitively motivated) LMs will not tell the whole story of human sentence processing van Schijndel and Linzen (2021); Huang et al. (2024); Prasad and Linzen (2024); Wang et al. (2024). We believe that the combination of such more linguistically motivated perspectives and our analysis will enrich the findings (e.g., we can also obtain probabilities from internal layers of neural parsers). This study itself contributes to clarifying to what extent surprisal from modern LMs, which are seemingly behaving similarly to humans Hu et al. (2024), can explain online human sentence processing. More specifically, this study also opens up several new questions. For example, how can the best layer for cognitive modeling be automatically found from a number of LLM layers? What are the information-theoretical insights for cognitive modeling and surprisal theory?
Appendix A Cross-lingual experiments

### v2

@@ 6.4 When are earlier layers advantageous?
We further explore more general trends on when and why earlier layers’ suprisal aligns better with human reading data. Following Oh and Schuler (2023b), we analyze by-token squared residual errors from regression models predicting human data (Section˜3.2). We specifically use the largest data of DC (FPGD) and identify tokens where the use of the best internal layer, rather than the last layer, notably reduces errors. [^11] Oh and Schuler (2023b) reported a misalignment between MSEs (residual errors) and log-likelihood scores due to the Euclidean norm penalty adopted in the lme4 package. This did not arise in our analysis with statsmodels. To address this question, we fit a linear regression model to explain the decreases in squared residual errors observed in each data point $w_{t}$ with an LM $\theta$, with the following by-token linguistic properties as features: $\mathrm{model}(\theta)$, $\mathrm{length}(w_{t})$, $\mathrm{freq}(w_{t})$, $\mathrm{position}(w_{t})$, $\mathrm{POS}(w_{t})$}.
@@
Figure 5: Scaling effect between $\Delta$LL and parameter counts in MECO. The grey lines are results relying on the last layer’s $\Delta$LLs, and the red lines rely on the best internal layers’ $\Delta$LLs. In all the languages, the negative correlation between parameter size and $\Delta$LL is mitigated to some extent, and in four languages, the relationship turned out to be positive.
(a) Multilingual models
(b) Monolingual models
Figure 6: Scaling effects in MECO German part
6.5 Mutlilingual generallity
Our main experiments were limited to the English language — what about the advantage of internal surprisal in other languages? Following the experiments of de Varda and Marelli (2023), we also explored the multilingual generality of our results, using FPGD data in ten languages recorded in MECO Siegelman et al. (2022) Five variants of multilingual XGLMs Lin et al. (2022) are used. [^12] We excluded Dutch, Hebrew, and Norwegian because these are not supported by XGLMs. In addition, due to the unavailability of tuned-lens for XGLMs, only logit-lens was used. Figure 5 shows the relationship between parameter size and $\Delta$LL, where the red markers and lines are based on the best internal layers, and the gray ones are from the final layers. First, the best layer typically outperforms the last layer of the respective model (86%=43/50 of the settings), as the best $\Delta$LLs (red markers) are generally higher than the last layer’s $\Delta$LLs (grey markers) in Figure 5, reproducing the findings in Section˜5.1.
Second, the negative correlation score between parameter size and $\Delta$LL consistently increases in all the languages, which partially aligns with Section˜5.2. Nevertheless, the scaling effect is still negative in some languages. We suspect that these mixed results might be biased by multilingual LMs, which are reported to process every language within the English subspace in their middle layers Wendler et al. (2024), leading to biased probability estimates. As a case study to handle this concern, we compared the results of multilingual XGMs (left part of Figure 6) with those of German monolingual GPT-2s (right part of Figure 6; see Appendix B for model details). The negative scaling effects are flipped to be positive with the use of monolingual LMs, implying that the inclusion of monolingual LMs further enhances the advantage of the internal layer’s surprisal.
7 Discussion
We lastly discuss the implications of our findings, connections to existing studies, and future works.
7.1 Connection to context sensitivity
We propose a perspective linking our findings to the context sensitivity of humans and LMs during sentence processing. Earlier layers may better model human reading behavior because they are less contextualized, reflecting the human-like tendency to process sentences under working memory constraints.
@@ 7.1 Connection to context sensitivity
Figure 7: The markers correspond to all the internal layers of our targeted LMs, which are sorted by relative layer depth (x-axis). Two types of scores (y-axis) are plotted: (i) Pearson correlation coefficient between each layer’s surprisal vs. less-contextualized bigram surprisal (blue); and (ii) each layer’s surprisal vs. well-contextualized LLM surprisal (red). We used tuned-lens results.
@@
We also confirmed the by-layer context-sensitivity of surprisal by analyzing the correlation between (i) intermediate layers’ surprisal and less-contextualized bigram surprisal [^13] Bigram LM is trained on OpenWebText Gokaslan and Cohen (2019) with the KenLM toolkit Heafield (2011); and (ii) intermediate layers’ surprisal and more contextualized surprisal from the LLM with the lowest PPL (OPT-66B). Figure 7 shows the above two types of correlations for each model’s layer, with the x-axis as the relative layer depth. This shows that earlier layers correlate more strongly with bigram surprisal (Pearson correlation coefficient $r$ for this relationship is --0.92), while later layers align with more accurate, well-contextualized surprisal ($r$=0.95). These findings reinforce the idea that earlier layers exhibit limited context sensitivity, while later layers are more contextualized and better suited for modeling data like MAZE, which demands higher contextualization.
7.2 Connection to spill-over effects
There is another possible explanation about the alignment of earlier/later layers with SPR/MAZE that was discussed in Section˜6.4. Self-paced reading and eye-tracking measures often exhibit spillover effects, where the processing of one word influences subsequent words Rayner (1998). This suggests that the comprehension of a word extends beyond the immediate moment, and the associated reading times may only capture an early stage of processing. In contrast, the MAZE task Forster et al. (2009); Boyce and Levy (2023), which measures the time taken to select a plausible continuation from two candidates, mitigates spillover effects and is thus expected to reflect the full process of word processing. Our findings — early LM layers are more closely aligned with gaze durations and self-paced reading times, while later layers show a stronger alignment with MAZE — align with this perspective.
7.3 Connection to LM-brain alignment study
Brain imaging data (e.g., fMRI) and reading behavior data have complementary advantages. Generally speaking, the former has a high spatial resolution (in which part of the brain particular processing is performed), while the latter has a high temporal resolution (how long does the processing take). In the former context of fMRI modeling research, layer-wise LM-human alignments, similarly to us, have been attempted Toneva and Wehbe (2019); Schrimpf et al. (2021); Caucheteux et al. (2023) and suggested that different brain areas better align with different LM layers. Our study is orthogonal to these studies with more focus on the temporal alignment between reading-time and layer-time scales, which is concurrently explored in Hu et al. (2025).
One additional difference with the above-mentioned fMRI studies is that they typically trained linear regression models to predict brain activity directly using $d$-dimentional LM internal representations $\bm{h}\in\mathbb{R}^{d}$ as features, instead of using surprisal measures $-\log p(\mathrm{word}|\mathrm{context})\in\mathbb{R}_{\geq 0}$. Thus, their results are not directly comparable with existing surprisal-based studies and may rather suffer from a confounding factor of different $d$ for different LMs when precisely discussing LM-scaling effects.
7.4 Connection to early exits of neural models’ prediction
In the engineering context, predictions from internal layers of deep neural networks (i.e., early exits of the results from internal layers) are used to improve inference efficiency by avoiding their overthinking Graves (2016); Banino et al. (2021), which is also called adaptive computation time Kaya et al. (2019); Zhou et al. (2020). Such a technique has also recently been employed to enhance interpretability research to identify at which layer a particular prediction shapes nostalgebraist (2020); Dar et al. (2023); Belrose et al. (2023).
7.5 Limitations toward surprisal theory
Recent studies have raised several issues, orthogonal to our study, on surprisal-based cognitive modeling, for example, on tokenizations Nair and Resnik (2023); Giulianelli et al. (2024a); Oh and Schuler (2024), LM training scenario Oh and Schuler (2023a), as well as more refined indicators of word predictability Pimentel et al. (2023); Giulianelli et al. (2024b); Opedal et al. (2024); Meister et al. (2024). More critically, the reliance solely on surprisal obtained from LMs tends to underestimate the significant slowdown of sentence processing against syntactic amguity and grammatical violation Wilcox et al. (2021); van Schijndel and Linzen (2021); Huang et al. (2024); Wang et al. (2024). This study, as an initial foray, is focused on cognitive modeling on naturalistic reading corpora, but including such experiments with controlled materials will enrich our findings on internal surprisals.
Despite such concerns regarding surprisal theory, we would like to highlight that our core proposal — using probability from internal layers in cognitive modeling — is not limited to surprisal theory in human sentence processing modeling, but can contribute to any cognitive modeling work that uses probability-related measures and neural models. For example, our analysis can easily be extended to other metrics such as entropy, entropy reduction Hale (2016), or any new probability-based measures combining probabilities from different layers Hu et al. (2025). Furthermore, our idea to focus on internal layers may also be combined with linguistically-motivated neural models, such as deep neural incremental parsers Dyer et al. (2016); Yoshida et al. (2021), to potentially combine linguistic theories (e.g., theory of syntax) with surprisal theory. Thus, any limitations of surprisal theory itself do not undermine the core contributions of this work.
@@ 8 Conclusions
Recent cognitive modeling studies have demonstrated a worse fit of surprisal from larger LMs to human reading time. In this paper, we argue that these negative results stem from the exclusive focus on the final layers of LMs. Instead, we observe that those from internal layers comparably or even better fit with human behavior and neurophysiology data, suggesting that the cognitive plausibility of larger LMs has been underestimated. Furthermore, different human measurements align with different layers, implying an intriguing parallel between temporal dynamics in human sentence processing and LM internal layers.
@@
The experiments can be extended to more types of human measures, such as first fixation time, go-past reading time, eye regressions, total fixation time, EEG data other than the N400 potential Federmeier et al. (2007) as well as FMRI data Shain et al. (2022). As a first step, we have begun with the measures of SPR, FPGD, N400, and MAZE, as they are typically used in cognitive modeling. We also conducted an exploratory analysis with other human measures (e.g., P600, second-pass gaze duration) on the UCL corpus (Table 7 in Appendix). For such an extended scope, we cannot confirm the implied relationship between fast–slow human responses and early–late LM layers, and our findings appear to hold primarily for commonly-analyzed human measures of SPR, FPGD, N400, and MAZE. It is also a critical question in psycholinguistics whether the distinction between “fast” (gaze durations) and “slow” (N400) human measures adopted in this study is truly related to some cognitive costs or the time indeed humans took to process the word. Especially for EEG, the delay may perhaps be just due to technical properties regarding the temporal resolutions.
Acknowledgment
We appreciate our TACL action editor and reviewers. We also thank Ryo Yoshida for constructive feedback on the earlier version of this work. This work was supported by JSPS KAKENHI Grant Number JP24H00087; JST PRESTO Grant Number JPMJPR21C2; JSPS Grant-in-Aid for Early-Career Scientists Grant Number JP23K16938.
Appendix A Details on psychometric predictive power
Here, we explain the $\Delta$LL (PPP) score more formally. Recall that we quantify the predictive contribution of word-by-word surprisal to human cognitive responses using a log-likelihood-based score, $\Delta$LL. Let $\bm{w}=[w_{1},\dots,w_{n}]^{\top}$ denote a sequence of words, and let $\bm{y}=[y_{1},\dots,y_{n}]^{\top}$ be the corresponding word-by-word human cognitive costs, e.g., reading times.
For each word $w_{i}$, we compute a surprisal value $s(w_{i})\in\mathbb{R}$, and a vector of baseline linguistic features $b(w_{i})\in\mathbb{R}^{d}$. We consider two linear models: a full model using both surprisal and baseline features, and a reduced model using only baseline features.
In both cases, we estimate the coefficients using ordinary least squares (OLS). The full model solves:
$\hat{\phi}_{s},\hat{\bm{\phi}}_{b}=\arg\min_{\phi_{s},\bm{\phi}_{b}}\sum_{i=1}^{n}\left(y_{i}-\phi_{s}s(w_{i})-\bm{\phi}_{b}^{\top}b(w_{i})\right)^{2}.$
We assume homoscedastic Gaussian noise and evaluate the total log-likelihood of the fitted model using the maximum likelihood estimate of the variance:
$\text{LL}_{\text{full}}=-\frac{n}{2}\log(2\pi\hat{\sigma}_{\text{full}}^{2})-\frac{n}{2},\quad\text{where}\quad\hat{\sigma}_{\text{full}}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{\phi}_{s}s(w_{i})-\hat{\bm{\phi}}_{b}^{\top}b(w_{i})\right)^{2}.$
The reduced model is obtained by regressing $y$ on baseline features alone:
$\hat{\bm{\phi}}_{b}^{\prime}=\arg\min_{\bm{\phi}_{b}}\sum_{i=1}^{n}\left(y_{i}-\bm{\phi}_{b}^{\top}b(w_{i})\right)^{2},$
with corresponding log-likelihood:
$\text{LL}_{\text{baseline}}=-\frac{n}{2}\log(2\pi\hat{\sigma}_{\text{baseline}}^{2})-\frac{n}{2},\quad\text{where}\quad\hat{\sigma}_{\text{baseline}}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{\bm{\phi}}_{b}^{\prime\top}b(w_{i})\right)^{2}.$
We define the psychometric predictive power of surprisal as the difference in log-likelihood between the full and reduced models:
$\Delta\text{LL}=\text{LL}_{\text{full}}-\text{LL}_{\text{baseline}}=\frac{n}{2}\log\left(\frac{\hat{\sigma}_{\text{baseline}}^{2}}{\hat{\sigma}_{\text{full}}^{2}}\right).$
This $\Delta$LL value quantifies the contribution of surprisal to predicting human sentence processing data, with higher values indicating stronger predictive power.
Appendix B Models
Table 4 shows the exact models we used.
Appendix C Details on cross-lingual experiments